# ViT-Linearizer: Distilling Quadratic Knowledge into Linear-Time Vision Models

### v1

@@ Abstract
Vision Transformers (ViTs) have delivered remarkable progress through global self-attention, yet their quadratic complexity can become prohibitive for high-resolution inputs. In this work, we present ViT-Linearizer, a cross-architecture distillation framework that transfers rich ViT representations into a linear-time, recurrent-style model. Our approach leverages 1) activation matching, an intermediate constraint that encourages student to align its token-wise dependencies with those produced by the teacher, and 2) masked prediction, a contextual reconstruction objective that requires the student to predict the teacher’s representations for unseen (masked) tokens, to effectively distill the quadratic self-attention knowledge into the student while maintaining efficient complexity. Empirically, our method provides notable speedups particularly for high-resolution tasks, significantly addressing the hardware challenges in inference. Additionally, it also elevates Mamba-based architectures’ performance on standard vision benchmarks, achieving a competitive 84.3% top-1 accuracy on ImageNet with a base-sized model. Our results underscore the good potential of RNN-based solutions for large-scale visual tasks, bridging the gap between theoretical efficiency and real-world practice.
@@ 1 Introduction
Recent advances in training foundational Vision Transformers (ViTs) [13] have led to significant breakthroughs in visual representation learning, with numerous self-attention-based architectures now dominating the applications in visual understanding, generation, and multi-modal reasoning tasks [42, 21, 7, 39, 34, 38, 44, 29, 1]. As self-attention requires each token to compute its correspondence with every other positions in the sequence, ViT models are able to produce robust token-level (i.e., patch-level) representations even in weakly supervised or unsupervised training scenarios [13, 7, 33, 11, 38, 60, 45]. Nonetheless, this strong local representation capability comes with a quadratic computation complexity with respect to sequence length, raising a considerable challenge in inference efficiency and hardware capacities when learning from long contexts. It is noteworthy that this computational overhead was not particularly prohibitive in standard benchmarks featuring medium-resolution and coarse-granularity visual tasks that were historically focused on [12, 15, 32, 59, 6, 17]. However, as the demand for processing high-resolution, high-fidelity visual inputs rapidly grows, the question of how to efficiently leverage the quadratic knowledge learned by ViTs has become increasingly important.
As alternatives, RNN-fashion token mixers such as Mamba [18, 10], RWKV [35], and xLSTM [4] have recently been introduced in vision tasks, showcasing competitive prediction results and superior accuracy-computation trade-offs compared with ViTs [63, 47, 14, 2]. These recurrent vision models, whose computational cost and memory requirements scale linearly with sequence length, emerge as a potential solution to address the computational explosive issues posed by the quadratic complexity of self-attention when processing long contexts. However, unlike ViTs—into which substantial effort and resources have already been invested—the exploration of recurrent vision models has so far been restricted to relatively small data scales and modest model sizes. These limitations motivate us to develop a cross-architecture distillation approach that can effectively transfer the capabilities of ViTs into the linear-time recurrent models such as Vision Mamba [63] and Adventurer [47].
@@
We term this new approach ViT-Linearizer. Overall, ViT-Linearizer confers two primary benefits that serve as the core contributions of this work. First, ViT-Linearizer effectively transfers the knowledge learned by ViTs under their quadratic complexity costs into linear-time recurrent vision models with minimal compromise in predictive performance, providing a promising solution to mitigating ViTs’ considerable inference cost and elevated memory demands. In our experiments, we successfully distill the ViT-Base encoder from CLIP [38] into a linear-complexity Adventurer-Base model [47], which ultimately achieves a highly competitive fine-tuning accuracy of 84.3% on ImageNet. As shown in Figure 1, when processing high-resolution inputs, our distilled model offers significant inference speedup (e.g., 4.2 for Cityscapes [9] semantic segmentation) compared to its ViT teacher. Moreover, from the perspective of improving the practical performance of recurrent vision models, ViT-Linearizer produces a new state of the art for the Mamba-based architecture on various vision tasks. Remarkably, we successfully boost Adventurer-Large’s test accuracy on the standard ImageNet-1k classification benchmark from 83.4% to 85.0%, demonstrating that given robust supervision, linear-complexity recurrent vision models can also achieve prediction performance on par with ViTs.
@@ 2 Related Work
While most modern vision architectures rely on self-attention or convolutional designs, there is a rising interest in recurrent approaches as they combine the advantages of the global receptive field and linear or near-linear time and space complexity. One of primary architectural design simply takes the macro structure of ViT as a blueprint and replaces self-attention by recurrent modules [63, 46, 2, 14, 27, 54, 47], complemented by additional designs that enable the otherwise unidirectional token mixer to operate smoothly in typical bidirectional vision tasks [63, 47]. In addition, other designs adopt hierarchical or hybrid structures to further enrich the model’s expressive capacity, often by incorporating multi-scale or progressive token downsampling so as to capture both fine-grained details and broader contextual dependencies in a unified framework [30, 20]. These recurrent-style vision models have recently attracted significant attention due to their intrinsic efficiency benefits, and have advanced rapidly with broad applications in both multi-modal understanding [37, 58] and generation tasks [24, 62, 41].
@@
Cross-architecture distillation offers a particularly efficient way to transfer pretrained knowledge. Compared to training from scratch that typically requires considerable resources and time, this approach allows newly proposed architectures to quickly assess their practical performance across various tasks at an early exploration stage. Notably, in the early stages of ViT research, DeiT [42] demonstrated how a well-pretrained CNN could be distilled into a ViT to achieve competitive performance, sparking numerous follow-up studies on effectively applying Transformer architectures in vision. More relevantly, Transformer-to-RNN distillation has recently garnered significant attention, with numerous pioneering efforts already demonstrating its effectiveness and feasibility in language and vision-language tasks [5, 25, 57, 56, 28, 49, 31]. This work aims to explore how to advance self-attention linearization from the perspective of purely vision-driven tasks, proposing a cross-architecture distillation framework that transfers the strong representational capabilities of pretrained Transformers into efficient recurrent models, thereby bridging the gap between theoretical efficiency and practical performance in the visual domain.
@@ 3.3 Masked Prediction
Similar to the earlier observations in pretraining ViTs [21, 3, 51, 61, 36], we find that masked prediction, instead of aligning the entire feature map with the supervision information, makes a substantial difference for the representational capability of Mamba-based architectures. Specifically, we follow the standard asymmetric structure [3, 61, 36] to feed the teacher model with the entire image and the student model with masked input which is implemented by randomly replacing a portion of patch tokens with a learnable [mask] token. The teacher and student models finally produce outputs , . We discard the unmasked positions and compute the masked prediction loss by

### v2

@@ Abstract
Vision Transformers (ViTs) have delivered remarkable performence through global self-attention, yet their quadratic complexity can become prohibitive for high-resolution inputs. In this work, we present ViT-Linearizer, a cross-architecture distillation framework that transfers rich ViT representations into a linear-time, recurrent-style model. Our approach leverages 1) activation matching, an intermediate constraint that encourages a student to align its token-wise dependencies with those produced by the teacher, and 2) masked prediction, a contextual reconstruction objective that requires the student to predict the teacher’s representations for unseen (masked) tokens, to effectively distill the quadratic self-attention knowledge into the student while maintaining efficient complexity. Empirically, our method provides notable speedups particularly for high-resolution tasks, significantly addressing the hardware challenges in inference. Additionally, it also elevates Mamba-based architectures’ performance on standard vision benchmarks, achieving a competitive 84.3% top-1 accuracy on ImageNet with a base-sized model. Our results underscore the good potential of RNN-based solutions for large-scale visual tasks, bridging the gap between theoretical efficiency and real-world effectiveness.
@@ 1 Introduction
Recent advances in training foundational Vision Transformers (ViTs) [13] have led to significant breakthroughs in visual representation learning, with numerous self-attention-based architectures now dominating the applications in visual understanding, generation, and multi-modal reasoning tasks [45, 21, 7, 42, 34, 38, 47, 29, 1]. As self-attention requires each token to compute its correspondence with every other positions in the sequence, ViT models are able to produce robust token-level (i.e., patch-level) representations even in weakly supervised or unsupervised training scenarios [13, 7, 33, 11, 38, 64, 49, 48]. Nonetheless, this strong local representation capability comes with a quadratic computation complexity with respect to sequence length, raising a considerable challenge in inference efficiency and hardware capacities when learning from long contexts. It is noteworthy that this computational overhead was not particularly prohibitive in standard benchmarks featuring medium-resolution and coarse-granularity visual tasks that were historically focused on [12, 15, 32, 63, 6, 17]. However, as the demand for processing high-resolution, high-fidelity visual inputs rapidly grows, the question of how to efficiently leverage the quadratic knowledge learned by ViTs has become increasingly important.
As alternatives, RNN-fashion token mixers such as Mamba [18, 10], RWKV [35], and xLSTM [4] have recently been introduced in vision tasks, showcasing competitive prediction results and superior accuracy-computation trade-offs compared with ViTs [67, 51, 14, 2]. These recurrent vision models, whose computational cost and memory requirements scale linearly with sequence length, emerge as a potential solution to address the computationally explosive issues posed by the quadratic complexity of self-attention when processing long contexts. However, unlike ViTs—into which substantial effort and resources have already been invested—the exploration of recurrent vision models has so far been restricted to relatively small data scales and modest model sizes. These limitations motivate us to develop a cross-architecture distillation approach that can effectively transfer the capabilities of ViTs into the linear-time recurrent models such as Vision Mamba [67] and Adventurer [51].
@@
We term this new approach ViT-Linearizer. Overall, the ViT-Linearizer confers two primary benefits that serve as the core contributions of this work. First, the ViT-Linearizer effectively transfers the knowledge learned by ViTs under their quadratic complexity costs into linear-time recurrent vision models with minimal compromise in predictive performance, providing a promising solution to mitigating ViTs’ considerable inference cost and elevated memory demands. In our experiments, we successfully distill the ViT-Base encoder from CLIP [38] into a linear-complexity Adventurer-Base model [51], which ultimately achieves a highly competitive fine-tuning accuracy of 84.3% on ImageNet. As shown in Figure 1, when processing high-resolution inputs, our distilled model offers significant inference speedup (e.g., 4.2 for Cityscapes [9] semantic segmentation) compared to its ViT teacher. Moreover, from the perspective of improving the practical performance of recurrent vision models, the ViT-Linearizer produces a new state of the art for the Mamba-based architecture on various vision tasks. Remarkably, we successfully boost Adventurer-Large’s test accuracy on the standard ImageNet-1k classification benchmark from 83.4% to 85.0%, demonstrating that given robust supervision, linear-complexity recurrent vision models can also achieve prediction performance on par with ViTs.
@@ 2 Related Work
While most modern vision architectures rely on self-attention or convolutional designs, there is a rising interest in recurrent approaches as they combine the advantages of the global receptive field and linear or near-linear time and space complexity. One of the primary architectural designs simply takes the macro structure of ViT as a blueprint and replaces self-attention by recurrent modules [67, 50, 2, 14, 27, 58, 51, 41, 39], complemented by additional designs that enable the otherwise unidirectional token mixer to operate smoothly in typical bidirectional vision tasks [67, 51]. In addition, other designs adopt hierarchical or hybrid structures to further enrich the model’s expressive capacity, often by incorporating multi-scale or progressive token downsampling so as to capture both fine-grained details and broader contextual dependencies in a unified framework [30, 20]. These recurrent-style vision models have recently attracted significant attention due to their intrinsic efficiency benefits, and have advanced rapidly with broad applications in both multi-modal understanding [37, 62] and generation tasks [24, 66, 44].
@@
Cross-architecture distillation offers a particularly efficient way to transfer pretrained knowledge. Compared to training from scratch that typically requires considerable resources and time, this approach allows newly proposed architectures to quickly assess their practical performance across various tasks at an early exploration stage. Notably, in the early stages of ViT research, DeiT [45] demonstrated how a well-pretrained CNN could be distilled into a ViT to achieve competitive performance, sparking numerous follow-up studies on effectively applying Transformer architectures in vision. More relevantly, Transformer-to-RNN distillation has recently garnered significant attention, with numerous pioneering efforts already demonstrating its effectiveness and feasibility in language and vision-language tasks [5, 25, 61, 60, 28, 53, 31, 40]. This work aims to explore how to advance self-attention linearization from the perspective of purely vision-driven tasks, proposing a cross-architecture distillation framework that transfers the strong representational capabilities of pretrained Transformers into efficient recurrent models, thereby bridging the gap between theoretical efficiency and practical performance in the visual domain.
@@ 3.3 Masked Prediction
Similar to earlier observations in pretraining ViTs [21, 3, 55, 65, 36], we find that masked prediction, instead of aligning the entire feature map with the supervision information, makes a substantial difference for the representational capability of Mamba-based architectures. Specifically, we follow the standard asymmetric structure [3, 65, 36] to feed the teacher model with the entire image and the student model with masked input which is implemented by randomly replacing a portion of patch tokens with a learnable [mask] token. The teacher and student models finally produce outputs , . We discard the unmasked positions and compute the masked prediction loss by